an investor invested a total of 2300 in two mutual funds.one fundearned a 4% profit while the other earned a 2%. if the investors total profits were $56, how much were invested in each mutual fund?
Answer by Boreal(15235)

MFund1=x
MFund2=2300-x
interest on the first is .04x
interest on the second is .02(2300-x)
They add to 56
.04x+46-.02x=56
.02x=10
x=$500 in MFund1 @4%=$20
2300-x=$1800 in MFund2 @2%=$36
They add to $56.